OpenStudy (anonymous):

Simplify (7g2 + 9g + 5) + (7g2 – 18g – 10) Answer 14g2 + 9g – 5 14g2 – 9g – 5 14g2 – 9g + 5 14g2 + 9g + 5

OpenStudy (anonymous):

in this case we can take off bracket so we have 7g2 + 9g + 5 + 7g2 – 18g – 10 sum cofficients of \[g^2\] and cofficient of g example: \[\Large 2x^2+3x^2=(2+3)x^2=5x^2\] in your case you have \[g^2\] try to sum them...

OpenStudy (anonymous):

thanks for the help!! :)

OpenStudy (anonymous):

My pleasure....